# The Compilation and Validation of the Spectroscopic Redshift Catalogs for the DESI-COSMOS and DESI-XMMLSS Fields

### v2

@@ Abstract
Over several dedicated programs that include targets beyond the main cosmological samples, the Dark Energy Spectroscopic Instrument (DESI) collected spectra for 311,102 unique objects in two fields centered on the COSMOS and XMM-LSS fields. In this work, we develop spectroscopic redshift robustness criteria for those spectra, validate these criteria using visual inspection, and provide two custom Value-Added Catalogs with our redshift characterizations. With these criteria, we reliably classify 217,037 galaxies below $z<1.6$, 9,688 quasars and 35,505 stars. As a critical element in characterizing the selection function, we provide the description of 70 different algorithms that were used to select these targets from imaging data. To facilitate joint imaging/spectroscopic analyses, we provide row-matched photometry from the Dark Energy Camera, Hyper-Suprime Cam, and public COSMOS2020 photometric catalogs. Finally, we demonstrate example applications of these large catalogs to photometric redshift estimation, cluster finding, and completeness studies.
@@ 1 Introduction
In this work, we present a custom evaluation of all spectra within the COSMOS and XMM-LSS fields. We modify the redshift quality cuts beyond those that were informed by visual inspection for DESI galaxies (Lan et al., 2023) and quasars (Alexander et al., 2023). The modified quality cuts increase the redshift purity for the main samples while minimally reducing the completeness in these fields. We use these new criteria to develop an algorithm for determining reliable redshifts for the wealth of supplemental programs within these two fields. In total, over 32 deg2, we find robust spectroscopic classification of 217,037 galaxies below $z<1.6$, 9,688 quasars and 35,505 stars. In addition, several dedicated programs acquired spectra for potential Lyman-break and Lyman-alpha emission galaxy targets (Ruhlmann-Kleider et al., 2024; White et al., 2024) with full redshift performance to be reported in the future. We exclude these targets from our study, as the automated redshift classification is still in development.
@@ 4.1 Bright Galaxy Survey
A majority of BGS spectra (98.3%) pass the cuts defined in this section. Based on these classifications, there are 20,825 unique BGS targets spectroscopically confirmed as galaxies in the DESI-COSMOS field and 17,770 unique BGS targets spectroscopically confirmed as galaxies in the DESI-XMMLSS field. The redshift distribution of BGS galaxies with reliable redshift estimations in the DESI-COSMOS field is shown in the right panel of Figure 2. In comparing our robustness criteria to the ‘fiducial’ criteria in Hahn et al. (2023), we estimate the catastrophic failure rate by comparing the redshift estimates for objects with multiple observations. Using the fiducial criteria of ZWARN = 0 and $\Delta\chi^{2}>40$, the main BGS sample retains 98.9% of its objects and has an implied catastrophic failure rate of 1.4% (144/10064). Our robustness criteria retains 98.3% of objects, with an implied catastrophic failure rate of 0.72% (73/9974).
@@ 4.2 Luminous Red Galaxies
There are 11,192 LRG targets spectroscopically confirmed as galaxies in the DESI-COSMOS field and 9,448 LRG targets spectroscopically confirmed as galaxies in the DESI-XMMLSS field with robust redshifts. The redshift distribution of the LRG sample is shown in the right panel of Figure 5.
@@ 4.3 Emission Line Galaxies
With these cuts, 80.9% of the primary ELG sample is retained. In total, there are 27,475 ELG galaxies with reliable redshift estimates in the DESI-COSMOS field and 20,529 in the DESI-XMMLSS field. The redshift distribution of the ELG sample is shown in the bottom panel of Figure 7.
@@ 4.4 Quasars
The primary quasar population for the DESI-COSMOS field consists of 3,086 quasars while the DESI-XMMLSS field consists of 3,507 unique quasars. Using visual inspection of repeat observations, the combination of Redrock with these two afterburners, we report a catastrophic failure rate of $\sim 2.6\%$.
@@ 4.6.1 Secondary and Special Continuum-Characterized Samples
PRINCIPAL (LRG) 30086
LOWZ_TIER3 1857
LOWZ_FAINT 1803
@@
LOW_Z 750
BGS_BGS_FAINT 665
LOWZ_TIER2 642
@@
BGS_BGS_FAINT_HIP 264
UNWISE_BLUE 185
LOWZ_TIER1 145
UNWISE_GREEN 38
@@ 4.6.2 Secondary and Special Emission Line Characterized Samples
ELG_PRINCIPAL 56346
ELG (SPECIAL) 13672
FILLER_HIP 3742
FILLER_LOP 1676
HSC_HIZ_SNE 590
DESI_ELG_LOP 539
DESI_ELG_HIP 411
@@
DESI_DEEP_HIP 640
DESI_DEEP_LOP 455
DESI_DEEP_HIP2 118
@@ 4.7 Secondary and Special Quasar samples
WISE_VAR_QSO 1215
PSF_OUT(]textttSPECTYPE=‘QSO’) 867
DESI_QSO 368
ISM_CGM_QGP 53
QSO_RED 10
@@ 4.8 Secondary and Special Stellar samples
MWS_MWS_BROAD 191
MWS_FAINT_RED 64
MWS_MWS_MAIN_RED 64
MWS_CALIB 26
BACKUP_CALIB 26
FAINT_HPM 20
MWS_MWS_WD 16
@@ 5 Value-Added Catalog
The catalogs have two main components: the main spectroscopic redshift catalog and supplemental photometric measurements. The redshift catalogs consist of 311,102 unique objects which are row matched across the photometric extensions. Here we describe the data provided in the publicly available VACs for both the DESI-COSMOS and DESI-XMMLSS fields.
@@ 5.1 Redshift Catalog
The main data product of the spectroscopic catalog is found in the SPECZ_CAT extension. This extension models the DESI EDR spectroscopic redshift catalog data model [^9] https://desidatamodel.readthedocs.io/en/latest/index.html including target location, identification and targeting information. These columns include: DESINAME, TARGET_RA, TARGET_DEC, PMRA, PMDEC, OBJTYPE, DESI_TARGET, SCND_TARGET, MWS_TARGET, BGS_TARGET, SV1_DESI_TARGET, SV1_SCND_TARGET, SV1_MWS_ TARGET, SV1_BGS_TARGET, SV3_DESI_TARGET, SV3_SCND_TARGET, SV3_MWS_TARGET, SV3_BGS_TARGET, SURVEY, PROGRAM, CHI2, DELTACHI2, SPECTYPE, SUBTYPE, ZWARN, IS_QSO_MGII, IS_QSO_QN_NEW_RR, and TSNR2_BGS/LRG/ELG/QSO/LYA. Information indicating if the target has corresponding RVS measurements or photometry from the appended photometric catalogs is found in columns: HAS_RVS, HAS_DECALS_PHOT, HAS_DECAM_PHOT, HAS_HSC_UD_PHOT,HAS_HSC_ WIDE_PHOT, HAS_MERIAN_PHOT, HAS_COSMOS2020_PHOT.
@@ 5.1.2 Targets observed in multiple programs
For all parameters associated with Redrock and quasar afterburners, we simply report the values associated with the best quality spectrum. The SURVEY and PROGRAM columns indicate the epoch from which the best spectrum was observed. We also include two columns describing the total coadded exposure time of the best spectrum, TOTAL_COADD_EXPTIME, and number of epochs of observation for the target, N_EPOCHS. All targeting information from each epoch will be included in the corresponding target mask columns. The reported emission line flux, emission line uncertainty and stellar fit parameters from the RVS pipeline represent the error weighted mean and error on the weighted mean over all quality spectra.
@@ 6.2 Spectroscopic Redshift Comparison
Complementing the DESI-COSMOS catalog, the spectroscopic redshift compilation from the COSMOS collaboration provides 48,575 quality redshifts ($CL~\textrm{(confidence level)}\geq 95\%$) from spectroscopy of 97,929 unique objects in the central 4 deg2 region (Khostovan et al., 2025). In the same footprint, the DESI-COSMOS catalog reports 86,049 quality redshifts. For objects classified as galaxies, the catastrophic rate ($\Delta v<1000$ km s-1) from the DESI catalog is expected to be less than 0.5%. Matching across the central 4 deg2 region, we find the 27,935 objects in common between the two catalogs. In Figure 18, a comparison between spectroscopic redshifts yields a 99.3% agreement within 1000 km s-1.
@@
The zCOSMOS catalog has 20,640 galaxies with high quality redshifts unique to the compilation, of which 3,333 are galaxies with $z>1.6$ that cannot be classified by the current version of the DESI pipeline due to limitations in the galaxy template set. The recent zCOSMOS catalog also includes physical parameter estimates such as stellar mass and star formation rates. The DESI-COSMOS offers 58,114 redshifts for $z<1.6$ galaxies in the central 4 deg2 footprint that are distinct from the COSMOS compilation. The DESI-COSMOS catalog includes an additional 131,702 unique redshifts in an extended 16 deg2. Both the DESI-COSMOS and DESI-XMMLSS catalogs provide prominent emission line flux estimates for all objects, well-defined target selection algorithms, and row matched photometry from public DECam, HSC and COSMOS2020 catalogs.
@@ 7 Summary
This paper presents the DESI spectroscopic redshift catalogs for the DESI-COSMOS and DESI-XMMLSS fields. We developed robust spectroscopic redshift classification based on the spectroscopic properties of each target class in these two fields. In combination, the two catalogs present spectroscopic redshifts for more than 311,000 unique objects across two 16 deg2 fields with over 262,000 quality redshifts. The spectroscopic redshifts in this catalog are highly robust, yielding less than 1% catastrophic failure rates across various samples. These samples include, but are not limited to, blue-star-forming galaxies, passive red galaxies, active galactic nuclei, Type Ia supernovae hosts, low redshift dwarf galaxies and many other target selections. Lastly, we showed the utility of these heterogeneous samples through broad applications to completeness studies, spectroscopic and photometric redshift validation, and galaxy cluster confirmation.
@@ Appendix B Targeting Information
LRG 11702 11510 98.4%
ELG 35522 29468 83.0%
QSO 4868 4374 89.9%
BGS_ANY 21691 21326 98.3%
MWS_ANY 17905 17866 99.8%
STD_FAINT 1106 1101 99.5%
STD_BRIGHT 564 561 99.5%
STD_WD 92 89 96.7%
SCND_ANY 156243 122237 78.2%
@@
DARK_TOO_HIP 150802 117677 78.0%
LOW_Z_TIER3 1915 1864 97.3%
WISE_VAR_QSO 1593 1482 93.0%
PSF_OUT_DARK 1405 1307 93.0%
LOW_Z 886 851 96.0%
PV_DARK_HIGH 813 520 64.0%
PV_BRIGHT_HIGH 736 443 60.2%
@@
BRIGHT_TOO_HIP 723 596 82.4%
LOW_Z_TIER2 706 665 94.2%
PSF_OUT_BRIGHT 645 622 96.4%
MWS_FAINT_BLUE 612 606 99.0%
UNWISE_BLUE 306 301 98.4%
PV_DARK_MEDIUM 208 199 95.7%
PV_BRIGHT_MEDIUM 201 198 98.5%
LOW_Z_TIER1 168 146 86.9%
PV_DARK_LOW 123 117 95.1%
PV_BRIGHT_LOW 97 90 92.8%
MWS_FAINT_RED 64 64 100.0%
@@
UNWISE_GREEN 56 55 98.2%
QSO_RED 27 23 85.2%
MWS_CALIB 26 26 100.0%
BACKUP_CALIB 26 26 100.0%
FAINT_HPM 25 20 80.0%
SN_HOSTS 21 21 100.0%
@@
MWS_CLUS_GAL_DEEP 17 14 82.4%
MWS_RR_LYRAE 6 6 100.0%
WD_BINARIES_BRIGHT 6 6 100.0%
@@
ELG_PRINCIPAL 81583 60766 74.5%
PRINCIPAL 33083 32278 97.6%
ELG 19007 13908 73.2%
FILLER_HIP 5270 4068 77.2%
LOWZ_FAINT 2808 1823 64.9%
BGS_BGS_BRIGHT 1809 1784 98.6%
FILLER_LOP 1757 1684 95.8%
DESI_LRG 1115 1109 99.5%
MERIAN_MS2 1026 539 52.5%
MERIAN_MS1 1004 858 85.5%
DESI_DEEP_HIP 932 656 70.4%
BGS_BGS_FAINT 689 683 99.1%
DESI_ELG_LOP 666 554 83.2%
DESI_QSO 627 572 91.2%
LOWZ_BRIGHT 593 569 96.0%
DESI_DEEP_LOP 577 484 83.9%
DESI_ELG_HIP 479 422 88.1%
BGS_BGS_FAINT_HIP 267 266 99.6%
MWS_MWS_MAIN_BLUE 266 265 99.6%
MWS_MWS_BROAD 192 191 99.5%
@@
LRG 9986 9738 97.5%
ELG 28058 22429 79.9%
QSO 5253 4714 89.7%
BGS_ANY 18359 18023 98.2%
MWS_ANY 10175 10141 99.7%
SCND_ANY 11222 9525 84.9%
STD_FAINT 820 818 99.8%
STD_BRIGHT 465 463 99.6%
STD_WD 70 69 98.6%
QSO_Z5 25 25 100.0%
@@
DARK_TOO_HIP 3364 2422 72.0%
PSF_OUT_DARK 2469 2336 94.6%
BRIGHT_TOO_HIP 1591 1425 89.6%
@@
BGS_BGS_BRIGHT 2729 2719 99.6%
DESI_DEEP_LOP 1981 1605 81.0%
DESI_LRG 1891 1877 99.3%
DESI_ELG_LOP 1734 1445 83.3%

### v3

@@ Abstract
Over several dedicated programs that include targets beyond the main cosmological samples, the Dark Energy Spectroscopic Instrument (DESI) collected spectra for 304,970 unique objects in two fields centered on the COSMOS and XMM-LSS fields. In this work, we develop spectroscopic redshift robustness criteria for those spectra, validate these criteria using visual inspection, and provide two custom Value-Added Catalogs with our redshift characterizations. With these criteria, we reliably classify 212,935 galaxies below $z<1.6$, 9,713 quasars and 35,222 stars. As a critical element in characterizing the selection function, we provide the description of 70 different algorithms that were used to select these targets from imaging data. To facilitate joint imaging/spectroscopic analyses, we provide row-matched photometry from the Dark Energy Camera, Hyper-Suprime Cam, and public COSMOS2020 photometric catalogs. Finally, we demonstrate example applications of these large catalogs to photometric redshift estimation, cluster finding, and completeness studies.
@@ 1 Introduction
In this work, we present a custom evaluation of all spectra within the COSMOS and XMM-LSS fields. We modify the redshift quality cuts beyond those that were informed by visual inspection for DESI galaxies (Lan et al., 2023) and quasars (Alexander et al., 2023). The modified quality cuts increase the redshift purity for the main samples while minimally reducing the completeness in these fields. We use these new criteria to develop an algorithm for determining reliable redshifts for the wealth of supplemental programs within these two fields. In total, over 32 deg2, we find robust spectroscopic classification of 212,935 galaxies below $z<1.6$, 9,713 quasars and 35,222 stars. In addition, several dedicated programs acquired spectra for potential Lyman-break and Lyman-alpha emission galaxy targets (Ruhlmann-Kleider et al., 2024; White et al., 2024) with full redshift performance to be reported in the future. We exclude these targets from our study, as the automated redshift classification is still in development.
@@ 4.1 Bright Galaxy Survey
A majority of BGS spectra (98.3%) pass the cuts defined in this section. Based on these classifications, there are 21,684 unique BGS targets spectroscopically confirmed as galaxies in the DESI-COSMOS field and 18,021 unique BGS targets spectroscopically confirmed as galaxies in the DESI-XMMLSS field. The redshift distribution of BGS galaxies with reliable redshift estimations in the DESI-COSMOS field is shown in the right panel of Figure 2. In comparing our robustness criteria to the ‘fiducial’ criteria in Hahn et al. (2023), we estimate the catastrophic failure rate by comparing the redshift estimates for objects with multiple observations. Using the fiducial criteria of ZWARN = 0 and $\Delta\chi^{2}>40$, the main BGS sample retains 98.9% of its objects and has an implied catastrophic failure rate of 1.4% (144/10064). Our robustness criteria retains 98.3% of objects, with an implied catastrophic failure rate of 0.72% (73/9974).
@@ 4.2 Luminous Red Galaxies
There are 11,565 LRG targets spectroscopically confirmed as galaxies in the DESI-COSMOS field and 9,736 LRG targets spectroscopically confirmed as galaxies in the DESI-XMMLSS field with robust redshifts. The redshift distribution of the LRG sample is shown in the right panel of Figure 5.
@@ 4.3 Emission Line Galaxies
With these cuts, 80.9% of the primary ELG sample is retained. In total, there are 30,264 ELG galaxies with reliable redshift estimates in the DESI-COSMOS field and 22,427 in the DESI-XMMLSS field. The redshift distribution of the ELG sample is shown in the bottom panel of Figure 7.
@@ 4.4 Quasars
The primary quasar population for the DESI-COSMOS field consists of 4,471 quasars while the DESI-XMMLSS field consists of 4,711 unique quasars. Using visual inspection of repeat observations, the combination of Redrock with these two afterburners, we report a catastrophic failure rate of $\sim 2.6\%$.
@@ 4.6.1 Secondary and Special Continuum-Characterized Samples
PRINCIPAL (LRG) 32124
LOWZ_TIER3 1875
LOWZ_FAINT 1819
@@
LOW_Z 842
BGS_BGS_FAINT 679
LOWZ_TIER2 655
@@
BGS_BGS_FAINT_HIP 263
UNWISE_BLUE 296
LOWZ_TIER1 141
UNWISE_GREEN 53
@@ 4.6.2 Secondary and Special Emission Line Characterized Samples
ELG_PRINCIPAL 58057
ELG (SPECIAL) 13671
FILLER_HIP 4051
FILLER_LOP 1679
HSC_HIZ_SNE 625
DESI_ELG_LOP 547
DESI_ELG_HIP 420
@@
DESI_DEEP_HIP 657
DESI_DEEP_LOP 484
DESI_DEEP_HIP2 132
@@ 4.7 Secondary and Special Quasar samples
WISE_VAR_QSO 1494
PSF_OUT(SPECTYPE=‘QSO’) 867
DESI_QSO 596
ISM_CGM_QGP 55
QSO_RED 22
LOW_MASS_AGN 3
@@ 4.8 Secondary and Special Stellar samples
MWS_MWS_BROAD 188
MWS_MWS_MAIN_RED 65
MWS_FAINT_BLUE 59
MWS_CALIB 24
BACKUP_CALIB 24
MWS_FAINT_RED 18
FAINT_HPM 18
MWS_CLUS_GAL_DEEP 15
MWS_MWS_WD 15
@@ 5 Value-Added Catalog
The catalogs have two main components: the main spectroscopic redshift catalog and supplemental photometric measurements. The redshift catalogs consist of 304,970 unique objects which are row matched across the photometric extensions. Here we describe the data provided in the publicly available VACs for both the DESI-COSMOS and DESI-XMMLSS fields.
@@ 5.1 Redshift Catalog
[…]
@@ 5.1.2 Targets observed in multiple programs
For all parameters associated with Redrock and quasar afterburners, we simply report the values associated with the best quality spectrum. The SURVEY and PROGRAM columns indicate the epoch from which the best spectrum was observed. We also include two columns describing the total coadded exposure time of the best spectrum, TOTAL_COADD_EXPTIME, and number of observations for the target, TOTAL_NUM_COADDS. All targeting information from each epoch will be included in the corresponding target mask columns. The reported emission line flux, emission line uncertainty and stellar fit parameters from the RVS pipeline represent the error weighted mean and error on the weighted mean over all quality spectra.
@@ 6.2 Spectroscopic Redshift Comparison
Complementing the DESI-COSMOS catalog, the spectroscopic redshift compilation from the COSMOS collaboration provides 48,575 quality redshifts ($CL~\textrm{(confidence level)}\geq 95\%$) from spectroscopy of 97,929 unique objects in the central 4 deg2 region (Khostovan et al., 2025). In the same footprint, the DESI-COSMOS catalog reports 83,925 quality redshifts. For objects classified as galaxies, the catastrophic rate ($\Delta v<1000$ km s-1) from the DESI catalog is expected to be less than 0.5%. Matching across the central 4 deg2 region, we find the 28,030 objects in common between the two catalogs. In Figure 18, a comparison between spectroscopic redshifts yields a 99.2% agreement within 1000 km s-1.
@@
The zCOSMOS catalog has 20,173 galaxies with high quality redshifts unique to the compilation, of which 3,333 are galaxies with $z>1.6$ that cannot be classified by the current version of the DESI pipeline due to limitations in the galaxy template set. The recent zCOSMOS catalog also includes physical parameter estimates such as stellar mass and star formation rates. The DESI-COSMOS offers 45,196 quality redshifts for $z<1.6$ galaxies in the central 4 deg2 footprint that are distinct from the COSMOS compilation. The DESI-COSMOS catalog includes an additional 128,630 unique redshifts in an extended 16 deg2. Both the DESI-COSMOS and DESI-XMMLSS catalogs provide prominent emission line flux estimates for all objects, well-defined target selection algorithms, and row matched photometry from public DECam, HSC and COSMOS2020 catalogs.
@@ 7 Summary
This paper presents the DESI spectroscopic redshift catalogs for the DESI-COSMOS and DESI-XMMLSS fields. We developed robust spectroscopic redshift classification based on the spectroscopic properties of each target class in these two fields. In combination, the two catalogs present spectroscopic redshifts for more than 304,000 unique objects across two 16 deg2 fields with over 257,000 quality redshifts. The spectroscopic redshifts in this catalog are highly robust, yielding less than 1% catastrophic failure rates across various samples. These samples include, but are not limited to, blue-star-forming galaxies, passive red galaxies, active galactic nuclei, Type Ia supernovae hosts, low redshift dwarf galaxies and many other target selections. Lastly, we showed the utility of these heterogeneous samples through broad applications to completeness studies, spectroscopic and photometric redshift validation, and galaxy cluster confirmation.
@@ Appendix B Targeting Information
LRG 11720 11565 98.7%
ELG 36128 30264 83.8%
QSO 4873 4471 91.8%
BGS_ANY 21865 21684 99.2%
MWS_ANY 19424 19401 99.9%
STD_FAINT 1191 1189 99.8%
STD_BRIGHT 618 618 100.0%
STD_WD 88 87 98.9%
SCND_ANY 146452 113971 77.8%
@@
DARK_TOO_HIP 139840 108331 77.5%
LOW_Z_TIER3 1909 1875 98.2%
WISE_VAR_QSO 1581 1494 94.5%
PSF_OUT_DARK 1414 1342 94.9%
LOW_Z 856 842 98.4%
PV_DARK_HIGH 816 528 64.7%
PV_BRIGHT_HIGH 809 519 64.2%
@@
LOW_Z_TIER2 686 655 95.5%
PSF_OUT_BRIGHT 651 640 98.3%
UNWISE_BLUE 298 296 99.3%
PV_DARK_MEDIUM 208 201 96.6%
PV_BRIGHT_MEDIUM 201 200 99.5%
LOW_Z_TIER1 162 141 87.0%
PV_DARK_LOW 130 126 96.9%
PV_BRIGHT_LOW 129 124 96.1%
BRIGHT_TOO_HIP 140 35 25.0%
MWS_FAINT_BLUE 61 59 96.7%
@@
UNWISE_GREEN 54 53 98.1%
QSO_RED 26 22 84.6%
MWS_CALIB 24 24 100.0%
BACKUP_CALIB 24 24 100.0%
FAINT_HPM 23 18 78.3%
SN_HOSTS 22 22 100.0%
@@
MWS_CLUS_GAL_DEEP 18 15 83.3%
MWS_FAINT_RED 18 18 100.0%
@@
ELG_PRINCIPAL 78072 58057 74.4%
PRINCIPAL 32881 32124 97.7%
ELG 18603 13671 73.5%
FILLER_HIP 5240 4051 77.3%
LOWZ_FAINT 2760 1819 65.9%
BGS_BGS_BRIGHT 1754 1752 99.9%
FILLER_LOP 1752 1679 95.8%
DESI_LRG 1114 1112 99.8%
MERIAN_MS2 984 530 53.9%
MERIAN_MS1 938 805 85.8%
DESI_DEEP_HIP 932 657 70.5%
BGS_BGS_FAINT 680 679 99.9%
DESI_ELG_LOP 645 547 84.8%
DESI_QSO 627 596 95.1%
LOWZ_BRIGHT 589 569 96.6%
DESI_DEEP_LOP 576 484 84.0%
DESI_ELG_HIP 470 420 89.4%
MWS_MWS_MAIN_BLUE 265 265 100.0%
BGS_BGS_FAINT_HIP 263 263 100.0%
FILLER_LOP 1752 1679 95.8%
MWS_MWS_BROAD 188 188 100.0%
@@
LRG 9985 9736 97.5%
ELG 28056 22427 79.9%
QSO 5251 4711 89.7%
BGS_ANY 18357 18021 98.2%
MWS_ANY 10092 10057 99.7%
STD_FAINT 770 768 99.7%
STD_BRIGHT 425 423 99.5%
STD_WD 66 64 97.0%
SCND_ANY 10926 9232 84.5%
@@
DARK_TOO_HIP 3345 2404 71.9%
PSF_OUT_DARK 2468 2335 94.6%
WISE_VAR_QSO 1856 1741 93.8%
@@
BGS_BGS_BRIGHT 2662 2652 99.6%
DESI_DEEP_LOP 1979 1603 81.0%
DESI_LRG 1890 1875 99.2%
DESI_ELG_LOP 1733 1444 83.3%